OpenStudy (anonymous):

Three friends went to dinner and paid a total of $57. Sarah paid $17 and Jenny and Bill split the rest of hte bill evenly. Which of the following equations best represents this situation? $17 + 2x = $57 3x = $40 $17 – 2x = $57 $57 + 3x = $17

OpenStudy (anonymous):

no ;p;

OpenStudy (anonymous):

hellp

OpenStudy (anonymous):

3x=40?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

40

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

17

OpenStudy (anonymous):

okayyy but i am in a hurry

OpenStudy (anonymous):

Now, can you plug these numbers in? Sarah + Jenny + Bill = total

OpenStudy (anonymous):

40

OpenStudy (anonymous):

17 20 20

OpenStudy (anonymous):

I'm sorry that you have no choice except follow MY way ;) Sarah + Jenny + Bill = total $17 + x + x = $57

OpenStudy (anonymous):

Do you still need further help?

OpenStudy (anonymous):

17-2x ?

OpenStudy (anonymous):

where do you get -2x?

OpenStudy (anonymous):

can u just tell me if its wrong or right?

OpenStudy (anonymous):

You tell me + x + x = - 2x ???? :0

OpenStudy (anonymous):

idk!

OpenStudy (anonymous):

im in a hurry...

OpenStudy (anonymous):

x + x = 2x like candy + candy = 2 candies ( common sense)

OpenStudy (anonymous):

ur to high and mighty to cut someone some slack...

OpenStudy (anonymous):

yeah unfortunetley ur not my teacher and its nun of ur biznesss weather i learn or not....

OpenStudy (anonymous):

Okie, don't forget that you're the one who post the question!

OpenStudy (anonymous):

and i wasnt looking for the u teach me i was looking for the answer